# A Theory of Semantic Communication

### v3

@@ III Problem Formulation
The word and syntax of a language determine the set of all possible messages, denoted by $\mathcal{S}$. Suppose $\mathcal{S}$ is finite or countably infinite, we define $\mathcal{S}\triangleq\{s_{m}:m\in[M]\}$, where $s_{m}$ denotes a message, $M$ the number of all possible messages, and we define $[M]\triangleq\{1,2,\ldots,M\}$.
@@
Consider the communication between a transmitter and a receiver. Suppose the transmitted message is $s\in\mathcal{S}$ and the received message is $\hat{s}\in\mathcal{S}$ [^1] More generally, one can define another message set $\hat{\mathcal{S}}$ at the receiver and define $\hat{s}\in\hat{\mathcal{S}}$. We characterize the technical channel by the transition probabilities from $s$ to $\hat{s}$ and denote it by $\left\{c(\hat{s}|s)\in[0,1]:s,\hat{s}\in\mathcal{S},\sum_{\hat{s}}c(\hat{s}|s)=1\right\}$, or the matrix form $\bm{C}\in I\!\!R^{M\times M}$. Given the process of expression and interpretation within the semantic language, the semantic channel can be written as
$\Pr(\hat{w}|w)=\sum_{s,\hat{s}}p(s|w)c(\hat{s}|s)q(w|\hat{s}).$ (5)
@@ III-A Semantic encoding
In principle, we can impose constraints on $u(s|w)$ such that the semantic encoding scheme complies with the semantic language. For example, if the semantic encoding scheme must comply with the interpretation, we can impose the constraint “$u(s|w)=0$ if $p(s|w)=0$”; if the transmitter is not allowed to deceive the receiver, we can impose the constraint “$u(s|w)=0$ if $q(w|s)=0$”; if the transmitter is not allowed to misdirect the receiver, we can impose the constraint “$u(s|w)\geq u(s^{\prime}|w)$ if $q(w|s)\geq q(w|s^{\prime})$”. In this paper, however, we consider a general setup and do not impose any constraints on $u(s|w)$.
@@
Without loss of generality, we assume $\mathcal{S}$ is an ordered set throughout the paper such that for any two messages $s_{m_{1}}$,$s_{m_{2}}$, $m_{1},m_{2}\in[M]$, we have $\ell(s_{m_{1}})\leq\ell(s_{m_{2}})$ if $m_{1}<m_{2}$.
@@
Given the above definitions, we are ready to describe a point-to-point semantic communication process with semantic encoding at the transmitter. As shown in Fig. 5(a), the transmitter first generates a meaning $w\sim p(w)$. To convey $w$ to the receiver, the transmitter constructs a message $s\sim u(s|w)$ following its semantic encoding scheme. After passing through the semantic channel, the receiver observes $\hat{s}\sim c(\hat{s}|s)$ and reconstructs a meaning $\hat{w}\sim q(\hat{w}|\hat{s})$ following the interpretation of the language. The main problem of semantic encoding is to characterize the optimal encoding scheme that achieves the best trade-off between the semantic distortion and semantic cost.
@@ IV-A Semantic distortion-cost region and function
$\underline{\text{P1}}:\left(L_{\text{min}},\min_{(L_{\text{min}},D)\in R_{\text{enc}}}D\right),\leavevmode\nobreak\ \underline{\text{P2}}:\left(\min_{(L,D_{\text{min}})\in R_{\text{enc}}}L,D_{\text{min}}\right),$
$\underline{\text{P3}}:\left(\max_{(L,D_{\text{min}})\in R_{\text{enc}}}L,D_{\text{min}}\right),\leavevmode\nobreak\ \underline{\text{P4}}:\left(L_{\text{max}},\min_{(L_{\text{max}},D)\in R_{\text{enc}}}D\right),$
$\overline{\text{P1}}:\left(L_{\text{min}},\max_{(L_{\text{min}},D)\in R_{\text{enc}}}D\right),\leavevmode\nobreak\ \overline{\text{P2}}:\left(\min_{(L,D_{\text{max}})\in R_{\text{enc}}}L,D_{\text{max}}\right),$
$\overline{\text{P3}}:\left(\max_{(L,D_{\text{max}})\in R_{\text{enc}}}L,D_{\text{max}}\right),\leavevmode\nobreak\ \overline{\text{P4}}:\left(L_{\text{max}},\max_{(L_{\text{max}},D)\in R_{\text{enc}}}D\right),$
@@
3:$\underline{\mathcal{S}}^{\prime}_{w}=\mathcal{S}$, $\underline{s}^{l}=s_{1}$; $\underline{\mathcal{S}}^{\prime\prime}_{w}=\mathcal{S}$, $\underline{s}^{r}=s_{M}$;
4:$\overline{\mathcal{S}}^{\prime}_{w}=\mathcal{S}$, $\overline{s}^{l}=s_{1}$; $\overline{\mathcal{S}}^{\prime\prime}_{w}=\mathcal{S}$, $\overline{s}^{r}=s_{M}$;
5:for $m=2,3,...,M$ do
@@
7:         $\underline{\mathcal{S}}^{\prime}_{w}=\underline{\mathcal{S}}^{\prime}_{w}\backslash s_{m}$;
@@
9:         $\underline{\mathcal{S}}^{\prime}_{w}=\underline{\mathcal{S}}^{\prime}_{w}\backslash\underline{s}^{l}$;
@@
14:         $\underline{\mathcal{S}}^{\prime\prime}_{w}=\underline{\mathcal{S}}^{\prime\prime}_{w}\backslash s_{M-m+1}$;
@@
16:         $\underline{\mathcal{S}}^{\prime\prime}_{w}=\underline{\mathcal{S}}^{\prime\prime}_{w}\backslash\underline{s}^{r}$;
@@
21:         $\overline{\mathcal{S}}^{\prime}_{w}=\overline{\mathcal{S}}^{\prime}_{w}\backslash s_{m}$;
@@
23:         $\overline{\mathcal{S}}^{\prime}_{w}=\overline{\mathcal{S}}^{\prime}_{w}\backslash\overline{s}^{l}$;
@@
28:         $\overline{\mathcal{S}}^{\prime\prime}_{w}=\overline{\mathcal{S}}^{\prime\prime}_{w}\backslash s_{M-m+1}$;
@@
30:         $\overline{\mathcal{S}}^{\prime\prime}_{w}=\overline{\mathcal{S}}^{\prime\prime}_{w}\backslash\overline{s}^{r}$;
@@
34:$\underline{\mathcal{S}}_{w}=\underline{\mathcal{S}}^{\prime}_{w}\cup\underline{\mathcal{S}}^{\prime\prime}_{w}$, $\overline{\mathcal{S}}_{w}=\overline{\mathcal{S}}^{\prime}_{w}\cup\overline{\mathcal{S}}^{\prime\prime}_{w}$.
@@ IV-B Characterizing the distortion-cost region
For a $w\in\mathcal{W}$, the six subsets are defined based on the properties of $s$. Take the subset $\underline{\mathcal{S}}^{\prime}_{w}$ for example. Each of $s$ in $\underline{\mathcal{S}}^{\prime}_{w}$ has the property that there exists no $s^{\prime}\in\mathcal{S}\backslash\{s\}$ that is more accurate and cost-efficient than $s$ at the same time. It will be shown later that, for each meaning $w$, we only need to consider the messages in $\underline{\mathcal{S}}_{w}$ and $\overline{\mathcal{S}}_{w}$ to construct the deterministic encoding schemes and characterize the distortion-cost region of semantic encoding. The procedures to discover the six sets for a $w$ are described in Algorithm 1. The properties of the six subsets are summarized in Appendix B.
@@
1:Input: $(\mathcal{W},\mathcal{S},\bm{P},\bm{Q})$, $\ell(s)$, $d(w,\widehat{w})$.
@@
3:$\left\{\overline{\bm{U}}^{(t)},\left(L_{\overline{\bm{U}}^{(t)}},D_{\overline{\bm{U}}^{(t)},\bm{Q}}\right):t=0,1,2,...,\overline{T}\right\}$.
@@ V-A Semantic decoding
$n^{\prime}_{m}$ $=$ $\operatorname*{arg\,min}_{n}\psi_{p}(w_{n},s_{m}),$ (23)
$n^{\prime\prime}_{m}$ $=$ $\operatorname*{arg\,max}_{n}\psi_{p}(w_{n},s_{m}).$ (24)
@@
To simplify notation, we denote by $\hat{w}_{q}(\hat{s})$ the meaning that a received message $\hat{s}$ will be mapped to under prior $q(w)$, i.e., $\hat{w}_{q}(\hat{s})=w_{n^{(q)}_{m}}$. The achieved distortion of $\bm{V}^{*}_{q}$ can be written as
$D_{\bm{P},\bm{V}^{*}_{q}}=\sum_{\hat{s}}\psi_{p}(\hat{w}_{q}(\hat{s}),\hat{s}).$ (27)
@@
Figure 7: (a) Probability space in which $\bm{\alpha}_{q}(\hat{s})$ resides is an equilateral triangle when $N=3$. (b) We partition the equilateral triangle into three quadrilaterals. For a received message $\hat{s}$, the Hamming distortion of semantic decoding is determined by whether $\bm{\alpha}_{p}(\hat{s})$ and $\bm{\alpha}_{q}(\hat{s})$ fall into the same quadrilateral.
A geometric interpretation of (30) is given as follows. First, we normalize $q(w)p(\hat{s}|w)$ to a distribution $\bm{\alpha}_{q}(\hat{s})$ with each element being
@@
The probability space in which $\bm{\alpha}_{q}(\hat{s})$ resides can be represented as a unit simplex. When $N=3$, for example, the probability space is an equilateral triangle, as shown in Fig. 7(a). Based on the semantic decoding rule, we partition the probability space into $N$ regions, corresponding to the $N$ meanings in $\mathcal{W}$. When $N=3$, the equilateral triangle can be partitioned into three quadrilaterals, as shown in Fig. 7(b). Under the prior $p(w)$ and $q(w)$, semantic decoding yields $M$ points $\left\{\bm{\alpha}_{p}(\hat{s}):\hat{s}\in\mathcal{S}\right\}$ and $\left\{\bm{\alpha}_{q}(\hat{s}):\hat{s}\in\mathcal{S}\right\}$, respectively.
In the above context, the geometric interpretation of (30) is that $\bm{\alpha}_{p}(\hat{s})$ and $\bm{\alpha}_{q}(\hat{s})$ fall into the same region of the probability space, $\forall\hat{s}\in\mathcal{S}$. That is, under the Hamming distortion, the optimality of semantic decoding is governed by a “partition distance” between $\bm{\alpha}_{p}(\hat{s})$ and $\bm{\alpha}_{q}(\hat{s})$, i.e., whether they fall into the same region, but not the conventional distance measure between two probability distributions, such as the L1 or Kullback-Leibler (KL) distance.
@@
$D_{\bm{P},\bm{Q}}-D_{\bm{P},\bm{V}^{*}_{q}}=$ (33)
$\sum_{\hat{s}}\left(\max_{w}p(w)p(\hat{s}|w)-\sum_{\hat{w}}p(\hat{w})p(\hat{s}|\hat{w})q(\hat{w}|\hat{s})\right)\geq 0.$
@@
Refinement 1: If $\exists\leavevmode\nobreak\ w^{\prime}\in\widehat{\mathcal{W}}(\hat{s})$, s.t.,
$\min_{w\in\mathcal{W}(\hat{s})}d(w,w^{\prime})\geq\max_{\begin{subarray}{c}w\in\mathcal{W}(\hat{s}),\\ \widetilde{w}\in\widehat{\mathcal{W}}(\hat{s})\backslash\{w^{\prime}\}\end{subarray}}d(w,\widetilde{w}),$ (36)
@@
Refinement 2: If $\exists\leavevmode\nobreak\ w^{\prime\prime}\in\widehat{\mathcal{W}}(\hat{s})$, s.t.,
$\max_{w\in\mathcal{W}(\hat{s})}d(w,w^{\prime\prime})\leq\min_{\begin{subarray}{c}w\in\mathcal{W}(\hat{s}),\\ \widetilde{w}\in\widehat{\mathcal{W}}(\hat{s})\backslash\{w^{\prime\prime}\}\end{subarray}}d(w,\widetilde{w}),$ (38)
@@
In the first refinement (37), there exists a $w^{\prime}$ that is strictly worse than any other $\widetilde{w}\in\widehat{\mathcal{W}}(\hat{s})$. This is possible because the interpretation of the language can be suboptimal for some distortion functions. In this case, we remove the mapping between $\hat{s}$ and $w^{\prime}$ and assign the probability $q(w^{\prime}|\hat{s})$ proportionally to other $\widetilde{w}\in\widehat{\mathcal{W}}(\hat{s})$. In the second refinement (39), there exists a $w^{\prime\prime}$ that is strictly better than any other $\widetilde{w}\in\widehat{\mathcal{W}}(\hat{s})$. As a result, we construct $v(w|\hat{s})$ such that $\hat{s}$ is deterministically decoded to $w^{\prime\prime}$ regardless of $q(w)$.
@@ V-B Combined semantic encoding and decoding (CSED)
[…]
In this example, the Hamming distortion can be reduced to 00 with either semantic encoding or semantic decoding. However, when both semantic encoding and decoding are performed, the Hamming distortion is $1$. This suggests that CSED does not necessarily improve upon separate semantic encoding or decoding.
@@
Following this idea, the receiver can further predict that the transmitter would improve its expression from $\bm{P}$ to $\bm{U}^{\prime}$ and optimize $\bm{Q}$ based on $\bm{U}^{\prime}$ instead of $\bm{Q}$, and so on. The resulting Hamming distortion would alternate between 00 and $1$, depending on how many times the transmitter and receiver have thought about each other’s encoding and decoding schemes. This phenomenon resembles human communication – two smart people may not be able to communicate efficiently as they would keep inferring each other’s expression and interpretation and changing theirs continuously.
@@ VI-A1 The message and meaning sets
There are two words ‘$\mathcal{U}$’ and ‘$\mathcal{R}$’ in the semantic language, representing the actions of the bug (i.e., “up” and “right”). A message is composed of a sequence of words, representing the sequential actions (i.e., a trajectory) of the bug starting from $(0,0)$. The message ‘$\mathcal{URR}$’, for example, means that the bug goes up, right, and right from $(0,0)$. The maximal number of moves in the grid world is $4$; hence, the number of possible messages is $\sum_{i=0}^{4}2^{i}=31$. The message set can be written as
$\mathcal{S}_{\text{all}}=\big{\{}\emptyset,\mathcal{U},\mathcal{R},\mathcal{UU},\mathcal{UR},\mathcal{RU},\mathcal{RR},\mathcal{UUU},\mathcal{UUR},\mathcal{URU},$
$\mathcal{URR},\mathcal{RUU},\mathcal{RUR},\mathcal{RRU},\mathcal{RRR},\mathcal{UUUU},\mathcal{UUUR},\mathcal{UURU},$
$\mathcal{UURR},\mathcal{URUU},\mathcal{URUR},\mathcal{URRU},\mathcal{URRR},\mathcal{RUUU},\mathcal{RUUR},$
$\mathcal{RURU},\mathcal{RURR},\mathcal{RRUU},\mathcal{RRUR},\mathcal{RRRU},\mathcal{RRRR}\big{\}}$
@@ VI-A2 Error-free technical communication
For a given message, the technical language maps the words ‘$\mathcal{U}$’ and ‘$\mathcal{R}$’ to bit sequences ‘00’ and ‘$10$’, respectively. The set of channel symbols can be constructed by substituting ‘$\mathcal{U}$’ and ‘$\mathcal{R}$’ in $\mathcal{S}_{\text{all}}$ with ‘00’ and ‘$10$’, respectively.

### v4

@@ III Problem Formulation
The word and syntax of a language determine the set of all possible messages, denoted by $\mathcal{S}$. Suppose $\mathcal{S}$ is finite or countably infinite, we define $\mathcal{S}\triangleq\{s_{m}:m\in[M]\}$, where $s_{m}$ denotes a message, $M$ the number of all possible messages, and we define $M]\triangleq\{1,2,\ldots,M\}$.
@@
Consider the communication between a transmitter and a receiver. Suppose the transmitted message is $s\in\mathcal{S}$ and the received message is $\hat{s}\in\mathcal{S}$. [^1] More generally, one can define another message set $\hat{\mathcal{S}}$ at the receiver and define $\hat{s}\in\hat{\mathcal{S}}$. We characterize the technical channel by the transition probabilities from $s$ to $\hat{s}$ and denote it by $\left\{c(\hat{s}|s)\in[0,1]:s,\hat{s}\in\mathcal{S},\sum_{\hat{s}}c(\hat{s}|s)=1\right\}$, or the matrix form $\bm{C}\in I\!\!R^{M\times M}$. Given the process of expression and interpretation within the semantic language, the semantic channel can be written as
[…]
@@ III-A Semantic encoding
In principle, we can impose constraints on $u(s|w)$ such that the semantic encoding scheme complies with the semantic language. For example, if the semantic encoding scheme must comply with the interpretation, we can impose the constraint “$u(s|w)=0$ if p(s|w)=0p(s|w)=0”; if the transmitter is not allowed to deceive the receiver, we can impose the constraint “$u(s|w)=0$ if q(w|s)=0q(w|s)=0”; if the transmitter is not allowed to misdirect the receiver, we can impose the constraint $u(s|w)\geq u(s^{\prime}|w)$ if $q(w|s)\geq q(w|s^{\prime})$”. In this paper, however, we consider a general setup and do not impose any constraints on $u(s|w)$.
@@
Without loss of generality, we assume $\mathcal{S}$ is an ordered set throughout the paper such that for any two messages $s_{m_{1}}$$s_{m_{2}}$, $m_{1},m_{2}\in[M]$, we have $\ell(s_{m_{1}})\leq\ell(s_{m_{2}})$ if $m_{1}<m_{2}$.
@@
[…]
@@ IV-A Semantic distortion-cost region and function
$\hskip-19.91684pt\underline{\text{P1}}:\left(L_{\text{min}},\min_{(L_{\text{min}},D)\in R_{\text{enc}}}D\right),\leavevmode\nobreak\ \underline{\text{P2}}:\left(\min_{(L,D_{\text{min}})\in R_{\text{enc}}}L,D_{\text{min}}\right),$
$\hskip-19.91684pt\underline{\text{P3}}:\left(\max_{(L,D_{\text{min}})\in R_{\text{enc}}}L,D_{\text{min}}\right),\leavevmode\nobreak\ \underline{\text{P4}}:\left(L_{\text{max}},\min_{(L_{\text{max}},D)\in R_{\text{enc}}}D\right),$
$\hskip-19.91684pt\overline{\text{P1}}:\left(L_{\text{min}},\max_{(L_{\text{min}},D)\in R_{\text{enc}}}D\right),\leavevmode\nobreak\ \overline{\text{P2}}:\left(\min_{(L,D_{\text{max}})\in R_{\text{enc}}}L,D_{\text{max}}\right),$
$\hskip-19.91684pt\overline{\text{P3}}:\left(\max_{(L,D_{\text{max}})\in R_{\text{enc}}}L,D_{\text{max}}\right),\leavevmode\nobreak\ \overline{\text{P4}}:\left(L_{\text{max}},\max_{(L_{\text{max}},D)\in R_{\text{enc}}}D\right),$
@@
[…]
5:for m=2,3,,Mm=2,3,...,M do
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ IV-B Characterizing the distortion-cost region
[…]
@@
1:Input: $(\mathcal{W},\mathcal{S},\bm{P},\bm{Q})$, $\ell(s)$, d(w,w^)d($\widehat{w}$).
@@
[…]
@@ V-A Semantic decoding
$n^{\prime}_{m}\hskip-5.69046pt$ $=$ $\hskip-5.69046pt\operatorname*{arg\,min}_{n}\psi_{p}(w_{n},s_{m}),$ (23)
$n^{\prime\prime}_{m}\hskip-5.69046pt$ $=$ $\hskip-5.69046pt\operatorname*{arg\,max}_{n}\psi_{p}(w_{n},s_{m}).$ (24)
@@
[…]
@@
[…]
@@
[…]
@@
$\hskip-14.22636ptD_{\bm{P},\bm{Q}}-D_{\bm{P},\bm{V}^{*}_{q}}=$ (33)
$\hskip-14.22636pt\sum_{\hat{s}}\left(\max_{w}p(w)p(\hat{s}|w)-\sum_{\hat{w}}p(\hat{w})p(\hat{s}|\hat{w})q(\hat{w}|\hat{s})\right)\geq 0.$
@@
[…]
@@
[…]
@@
[…]
@@ V-B Combined semantic encoding and decoding (CSED)
Suppose the meaning set is $\mathcal{W}=\{\text{yes},\text{no}\}$, the message set is $\mathcal{S}=\{\text{head nod},\text{head shake}\}$, and the costs of the two messages are the same. The expression and interpretation of the semantic language are quite opposite, as shown in Fig. 8. At the transmitter, “head nod” stands for “yes” and “head shake” stands for “no”, while at the receiver, “head nod” stands for “no” and “head shake” stands for “yes”. Suppose the prior distribution $p(w)$ is known at the receiver and we consider the Hamming distortion.
In this example, the Hamming distortion can be reduced to 0 with either semantic encoding or semantic decoding. However, when both semantic encoding and decoding are performed, the Hamming distortion is $1$. This suggests that CSED does not necessarily improve upon separate semantic encoding or decoding.
@@
Following this idea, the receiver can further predict that the transmitter would improve its expression from $\bm{P}$ to $\bm{U}^{\prime}$ and optimize $\bm{Q}$ based on $\bm{U}^{\prime}$ instead of $\bm{Q}$, and so on. The resulting Hamming distortion would alternate between 0 and $1$, depending on how many times the transmitter and receiver have thought about each other’s encoding and decoding schemes. This phenomenon resembles human communication – two smart people may not be able to communicate efficiently as they would keep inferring each other’s expression and interpretation and changing theirs continuously.
@@ VI-A1 The message and meaning sets
There are two words $\mathcal{U}$’ and $\mathcal{R}$’ in the semantic language, representing the actions of the bug (i.e., “up” and “right”). A message is composed of a sequence of words, representing the sequential actions (i.e., a trajectory) of the bug starting from $(0,0)$. The message $\mathcal{URR}$’, for example, means that the bug goes up, right, and right from $(0,0)$. The maximal number of moves in the grid world is $4$; hence, the number of possible messages is $\sum_{i=0}^{4}2^{i}=31$. The message set can be written as
$\hskip-18.49411pt\mathcal{S}_{\text{all}}=\big{\{}\emptyset,\mathcal{U},\mathcal{R},\mathcal{UU},\mathcal{UR},\mathcal{RU},\mathcal{RR},\mathcal{UUU},\mathcal{UUR},\mathcal{URU},$
$\hskip-18.49411pt\mathcal{URR},\mathcal{RUU},\mathcal{RUR},\mathcal{RRU},\mathcal{RRR},\mathcal{UUUU},\mathcal{UUUR},\mathcal{UURU},$
$\hskip-18.49411pt\mathcal{UURR},\mathcal{URUU},\mathcal{URUR},\mathcal{URRU},\mathcal{URRR},\mathcal{RUUU},\mathcal{RUUR},$
$\hskip-18.49411pt\mathcal{RURU},\mathcal{RURR},\mathcal{RRUU},\mathcal{RRUR},\mathcal{RRRU},\mathcal{RRRR}\big{\}}$
@@ VI-A2 Error-free technical communication
For a given message, the technical language maps the words $\mathcal{U}$’ and $\mathcal{R}$’ to bit sequences ‘0’ and ‘$10$’, respectively. The set of channel symbols can be constructed by substituting $\mathcal{U}$’ and $\mathcal{R}$’ in $\mathcal{S}_{\text{all}}$ with ‘0’ and ‘$10$’, respectively.